# Detecting Neurocognitive Disorders through Analyses of Topic Evolution and Cross-modal Consistency in Visual-Stimulated Narratives

### v1

@@ V-A Experimental Setup
In this study, we conduct both classification and regression tasks to evaluate the performance of various feature sets and classifier models. Eight systems are designed, each utilizing distinct feature sets. System 1 employs nine acoustic features, while System 2 incorporates 13 linguistic features both introduced in Section IV-A. System 4 utilizes features derived from reference-based analysis as described in Section IV-B, and System 5 employs statistical features extracted from the DTM model outlined in Section IV-C. Systems 3, 6, and 7 leverage different combinations of the feature sets used in the aforementioned systems. As these features are all statistical in nature, we first apply Principal Component Analysis (PCA) with 5 components to reduce dimensionality and then utilize a Support Vector Machine (SVM) for prediction, with a Radial Basis Function (RBF) kernel and soft margin $C=1$. The hyper-parameters “components” of PCA, and “kernel” and “soft margin” of SVM are grid searched using System 7 with all statistical features. System 8, the TITAN model, as depicted in Fig. 4, is connected with a fully-connected linear layer as the task head. To align with PCA in Systems 1 to 7, the projection dimension $H^{\prime}$ of TITAN is also set as 5.
@@ V-B Evaluation Protocol
For the evaluation, we employed multiple metrics across both classification and regression experiments. For the classification task, we used the F1-score and AUC as the primary metrics, supplemented by recall (Rec.), precision (Prec.), and accuracy (Acc.), reported on the positive class (NCD). Despite of AUC, the classification metrics were computed with a default threshold set at 0.5. For regression, we used the coefficient of determination ($R^{2}$, ranging from $\infty$- ∞ to 1) and Root Mean Square Error (RMSE, ranging from 0 to +$\infty$+ ∞) as metrics. To ensure robust evaluation, we averaged the results from the final five epochs of each experiment. Notably, both the ADReSS and CU-MARVEL-RABBIT test sets were unseen and maintained an equal distribution of positive and negative samples, ensuring a balanced evaluation.
@@ V-C Results
Another finding is that macrostructural statistics (Sys. 6) outperform microstructural statistics (Sys. 3), achieving F1 scores of 0.6804 versus 0.5843 in classification, and R2 of 0.2864 versus 0.1904, indicating the efficacy of discourse-level analysis on NCD narratives.
@@ VI Analysis & Discussion
Figure 5: Correlation coefficients between proposed features and NCD labels, highlighting the statistical significance of each feature on the x-axis.
@@ VI-A Comparison with previous literature
To evaluate the performance of the models, we conducted a cross-comparison study on the semi-public ADReSS dataset. The results in Table IV demonstrates that the proposed TITAN model (Sys.8) achieves superior performance with the F1 score of 0.8889, exceeding others by at least 3.13%.
@@ VI-B Statistical Features Correlation Ranking
To evaluate the significance of the proposed statistical features, we performed Spearman’s rank correlation test to rank all features based on the absolute values of correlation coefficients. As illustrated in Fig. 5, the correlation distribution indicates that the microstructural features demonstrate greater importance than macrostructural features. Notably, nine of the top ten features are derived from microstructural analysis, highlighting the effectiveness of these features. Specifically, DTM-based metrics such as “topic consistency,” “rate of change,” and several reference-based coverage metrics ranked relatively high with statistical significance ($p<0.001$), suggesting their potential as indicative features. Furthermore, many features associated with narrative macrostructural abilities display highly negative correlations with NCDs, whereas most pause-related features exhibit positive correlations with NCDs. These findings underscore the potential utility of these features for effective classification and a deeper understanding of NCD-related characteristics.
VI-C Text-Visual Alignment Differences between NCDs and HCs
Figure 6: Averaged correlation maps between image embeddings (y-axis) and text embeddings (x-axis) from HCs (a) and NCDs (b) using TITAN.
Fig. 6 quantifies the cross-modal alignments among individuals of HCs and NCDs. The highlighted cells in the figure represent regions with high text-visual correlation, indicating stronger alignment between the image embeddings (y-axis) and the text embeddings (x-axis). Notably, The HC narratives (Fig. 6a) exhibited consistently relatively high correlation scores across the heatmap diagonal, signifying cross-modal alignment throughout the narrative sequence. In contrast, the NCD narratives (Fig. 6b) displayed dispersed attention scores, reflecting disjointed narrative elements corresponding to the image stimuli. This distinction underscores TITAN’s ability of modeling cross-modal consistency for narrative analysis, as well as its potential in identifying and assessing cognitive decline.
VI-D Exemplar Study on NCD Detection Task
Figure 7: Example of an NCD narrative, where the Cantonese words are transcribed by ASR and English words are translated by GPT-4o.
To elaborate the investigated patterns in this work, we select a specific example with NCD from the unseen test set, as shown in Fig. 7. The Cantonese transcription is obtained using SenseVoice-Small, and the translation is provided by GPT-4o [81] to convey key meanings of it, which omits repetitions, dysfluencies and stop words. It is observable that this narrative misses many animals and events from the stimuli, such as squirrel, goose, and their related events. Additionally, there are few transitions between sentences or events during the narrating process.
To better understand these patterns, especially the proposed features, we quantify their contributions of classifying this example using the SHAP (SHapley Additive exPlanations) value [82], as shown in Fig. 8. SHAP values indicate contribution of each feature during the process of classification, showing how each feature influences the final decision starting from the classification boundary. Fig. 8 compares the contributions in the systems using microstructural and macrostructural features respectively. For the classifier using microstructural features, the top contributing feature “speech duration” tends to correctly predict the case, but most other features hinder this prediction. Conversely, the classifier using macrostructural features highlights the rate of change (topic evolution) and reference-based metrics as key indicators, which help make a correct prediction. As a result, the microstructural analysis fails to accurately detect the NCDs, whereas the macrostructural analysis successfully identifies them. These distinctions underscore the importance of macrostructural patterns in distinguishing NCDs.
@@ VI-E Ablation Study of TITAN
We conducted an ablation study to assess the contributions of different components of the proposed TITAN model, as detailed in Table V. By removing specific parts (denoted as ‘w/o’), we observed their impact on performance for the NCD detection task. The results show that each component is essential, with the absence of the image embedding (w/o Image emb) causing the most significant performance drop, highlighting its critical role in the model’s efficacy.
TABLE V: Ablation study of TITAN (Classification)
@@
(a) Microstructural (Ground-truth: 1, Prediction Probability: 0.3159)
(b) Macrostructural (Ground-truth: 1, Prediction Probability: 0.7059)
Figure 8: Feature contributions for the classifiers using (a) microstructural features and (b) macrostructural features. The x-axis and bars represent the magnitude (SHAP values) and direction (colors) of feature contributions. The y-axis lists the names and standardized values of the top features.
@@ VII Conclusion & Future Work
Macrostructural narrative analysis, which measures global coherence and cohesion at the narrative level, is a vital research scope for NCD detection. However, existing studies typically ignore its temporal aspects and the dynamic topic evolution in NCDs narratives. In this study, we proposed two dynamic macrostructural modeling approaches, DTM-based temporal analysis and the Text-Image Temporal Alignment Network (TITAN), to measure the temporal alignment between the narrative topics and the implicit topics in references or visual stimuli. The DTM-based temporal analysis revealed the erratic topic evolution of NCDs, while TITAN further extracted the referenced dynamic topics from the sequential images in a more fine-grained manner. These approaches were evaluated on the Cantonese CU-MARVEL-RABBIT corpus for classification and regression tasks. TITAN achieved the highest performances in both tasks, surpassing the baseline feature sets and DTM-based metrics. Additional cross-comparison on the English ADReSS corpus and regression task further validated the effectiveness of the dynamic macrostructural modeling approaches.
We anticipate the temporal and macrostructural approaches to significantly contribute to the early screening and detection of NCDs, providing a more comprehensive framework for analyzing narratives from these perspectives. While these approaches can still be technically optimized, the concepts underlying them offer new insights into quantifying language use that may indicate cognitive decline, and enhance the understanding of NCDs in visual-stimulated narratives. Moreover, they can be extended to enhance multimodal fusion of more modalities, support multilingual generation due to their language-agnostic nature, and detect other spoken disorders.

### v2

@@ V-A Experimental Setup
In this study, we conduct both classification and regression tasks to evaluate the performance of various feature sets and classifier models. Eight systems are designed, each utilizing distinct feature sets. System 1 employs 10 acoustic features, System 2 incorporates 13 linguistic features, and System 3 combines these microstructural features. System 4 utilizes 16 proposed reference-based metrics, and System 5 employs 6 proposed DTM-based features. System 7 combines these macrostructural features, while System 7 integrates all of these microstructural and macrostructural features. As these features are all statistical in nature, we first apply Principal Component Analysis (PCA) with 5 components to reduce dimensionality and then utilize a Support Vector Machine (SVM) for prediction, with a Radial Basis Function (RBF) kernel and soft margin $C=1$. The hyper-parameters “components” of PCA, and “kernel” and “soft margin” of SVM are grid searched using System 7 with all statistical features. System 8, the TITAN model, as depicted in Fig. 4, is connected with a fully-connected linear layer as the task head. To align with PCA in Systems 1 to 7, the projection dimension $H^{\prime}$ of TITAN is also set as 5.
@@ V-B Evaluation Protocol
For the evaluation, we employed multiple metrics across both classification and regression experiments. For the classification task, we used the F1-score and AUC as the primary metrics, supplemented by recall (Rec.), precision (Prec.), and accuracy (Acc.), reported on the positive class (NCD). Despite of AUC, the classification metrics were computed with a default threshold set at 0.5. For regression, we normalized the ground-truth labels to the range [0, 1] and used the coefficient of determination ($R^{2}$, ranging from $\infty$- ∞ to 1) and Root Mean Square Error (RMSE, ranging from 0 to +$\infty$+ ∞) as metrics. To ensure robust evaluation, we averaged the results from the final five epochs of each experiment. Notably, both the ADReSS and CU-MARVEL-RABBIT test sets were unseen and maintained an equal distribution of positive and negative samples, ensuring a balanced evaluation.
@@ V-C Results
Another finding is that macrostructural statistics (Sys. 6) outperform microstructural statistics (Sys. 3), achieving F1 scores of 0.6947 versus 0.5843 in classification, and R2 of 0.2233 versus 0.1604 in regression, indicating the efficacy of discourse-level analysis on NCD narratives.
@@ VI Analysis & Discussion
Figure 5: Correlation coefficients between proposed features and NCD labels, highlighting the statistical significance of each feature on the x-axis. The amplitudes of the green (macrostructural) and red (microstructural) bars represent correlation values, indicating that most of the top-ranked features are macrostructural.
@@ VI-A Comparison with previous literature
To evaluate the performance of the models, we conducted a cross-comparison study on the public ADReSS dataset. The results in Table IV demonstrates that the proposed TITAN model (Sys.8) achieves superior performance with the F1 score of 0.8889, exceeding others by at least 3.13% relatively.
@@ VI-B Statistical Features Correlation Ranking
To evaluate the significance of the proposed statistical features, we performed Spearman’s rank correlation test to rank all features based on the absolute values of correlation coefficients. As illustrated in Fig. 5, the correlation distribution indicates that the macrostructural features demonstrate greater importance than microstructural features. Notably, DTM-based metric “topic consistency”, acoustic feature “speaking rate”, and several reference-based coverage metrics ranked relatively high with statistical significance ($p<0.001$), suggesting their potential as indicative biomarkers. Furthermore, many features associated with narrative macrostructural abilities display highly negative correlations with NCDs, whereas most pause-related features exhibit positive correlations with NCDs. These findings underscore the potential utility of these features for effective classification and a deeper understanding of NCD-related characteristics.
VI-C Interpretable Feature Contribution Analysis
To elucidate the role of the proposed features during the model’s decision pathways, we quantify their contribution using SHapley Additive exPlanations (SHAP) values [85]. SHAP is an interpretable method rooted in cooperative game theory that attributes the prediction of a model to each input feature fairly and consistently using Shapley values. Specifically, for each feature $x_{i}\in X$ (the complete feature set), the SHAP value $\phi_{i}$ represents its average differences in the model’s predictions when the target feature is added to all possible subsets $S$ of other features, i.e.
$\phi_{i}=\sum_{S\subseteq X\backslash\{x_{i}\}}w_{i}{[f(S\cup\{x_{i}\})-f(S)]},$
where $w_{i}=|S|!(|X|-|S|-1)!/|X|!$ denotes the weight accounting for subset permutations, ||$\cdot$|| ⋅ | denotes set cardinality, and $f(S)=ln{\frac{P(\hat{y}|S)}{1-P(\hat{y}|S)}}$ is the log-odds ratio of the model’s output probability $P$ based on given inputs $S$.
Figure 6: Global SHAP value distribution (beeswarm plot) across all test samples in CU-MARVEL-RABBIT. Features are ranked vertically by mean absolute SHAP values (most impactful at top). Each point represents a single instance: horizontal position indicates the SHAP value (negative/positive impact), color reflects the standardized feature value (blue = low, red = high), and point density along each feature shows value distribution.
This formulation enables global feature importance analysis through aggregation of SHAP values across the entire unseen test set. Figure 6 presents the global feature contributions ranked by their absolute SHAP values. Notably, the proposed macrostructural features, including “topic_variability”, “topic_change_rate”, “topic_consistency” and “ROUGE-L”, emerged as the top-ranked global contributors. These findings highlight the critical role of macrostructural patterns in distinguishing NCDs from HCs. The consistent directionality of their SHAP values (positive/negative associations with NCD classification) further validates their reliable discriminative power across diverse instances. For example, the NCD narrative in Fig. 2(b) exhibits low topic variability (original value: 0.25, normalized value: -1.71), high topic change rate (original: 0.61, normalized: 0.36), and low topic consistency (original: 0.5, normalized: -1.42), contributing to the model’s prediction of NCD with a high probability (0.99).
VI-D Text-Visual Alignment Differences between NCDs and HCs
Figure 7: Relative differences in text-visual correlation maps between NCD and HC groups: Cool colors represent areas of lower correlation in the NCD group; warm colors represent areas of higher correlation in the NCD group. The text timeline is resampled (via undersampling or interpolation) to match the number of images for visualization.
Fig. 7 presents the relative differences in cross-modal alignments between individuals of HCs and NCDs. To match the varying lengths of text segments with the number of images, the text timeline is resampled (through under-sampling or interpolation) to a fixed length of 15 for visualization. The colormap highlights regions of text-visual correlation divergence, with cool colors indicating diminished alignment in NCD narratives and warm colors representing higher correlation for the NCD group. Notably, the NCD narratives exhibit relatively lower text-visual correlation across the heatmap diagonal, signifying their disrupted cross-modal consistencies in the VSN task. These findings demonstrate TITAN’s ability to model cross-modal consistency for narrative analysis, providing a new dimension for NCD detection.
@@ VI-E Ablation Study of TITAN
We conducted an ablation study to assess the contributions of different components (e.g., “RoPE”, “Project-Up&Down”, “Attention”, and “Image embeddings”) of the proposed TITAN model, as detailed in Table V. By removing each component (denoted as ‘w/o’), we observed their impacts on performance for the NCD detection task. The results show that each component is essential, with the absence of the image embedding (w/o Image emb) causing the most significant performance drop, highlighting its critical role in the model’s efficacy.
TABLE V: Ablation study of TITAN by removing each module (“w/o” means “without”) in the NCD detection task. The best results are marked in red, and the second best are underlined.
@@ VII Conclusion & Future Work
VSN serves as both a production and reflection of integrated cognitive abilities, offering a promising avenue to detect NCDs in an automatic and holistic way. While macrostructural narrative analysis is vital for measuring global coherence in NCDs, the patterns of topic evolution and cross-modal consistency that inherent in VSNs remain underexplored. To address this, we proposed two dynamic macrostructural modeling approaches, DTM-based temporal analysis and the TITAN model, to measure the temporal alignment between the narrative topics and the implicit topics in references or visual stimuli. The DTM-based temporal analysis captures the topic evolution of VSNs, while TITAN further models the referenced dynamic topics from the sequential images in a more fine-grained manner. These approaches were evaluated on the Cantonese CU-MARVEL-RABBIT corpus, with TITAN achieving superior performances in both classification and regression tasks. Additional cross-comparison on the English ADReSS corpus and feature contribution analysis further validated the effectiveness of the dynamic macrostructural modeling approaches.
We anticipate the temporal and macrostructural approaches (like TITAN and DTM-based models) to significantly contribute to the early screening and detection of NCDs, providing a more comprehensive framework for analyzing narratives from these perspectives. While these approaches can further be technically optimized, the concepts underlying them offer new insights into quantifying language use that may indicate cognitive decline, and enhance the understanding of NCDs in VSNs. Unlike approaches that employ detection models as black boxes, our proposed methods offer enhanced interpretability through topic evolution analysis and cross-modality correlation analysis, potentially providing clinicians with more transparent insights into the decision-making process. Moreover, they can be extended to enhance multimodal fusion of more modalities, support multilingual generation due to their language-agnostic nature, and detect other spoken disorders.